Among 157 African-American men, the mean systolic blood pressure was 146 mm Hg with a standard deviation of 27. We wish to know if on the basis of these data, we may conclude that the mean systolic blood pressure for a population of African-American is 140


1
Expert's answer
2021-04-15T07:52:30-0400

"n = 157 \\

\bar{X} = 146 \\

s = 27"

Assumption: population is not normal, σ2 is unknown, α=0.01

"H_0 : μ=140 \\

H_1 : μ≠140"

Test Statistic:

"Z = \frac{\bar{X}-μ}{s/ \sqrt{n}} \\

Z = \frac{146-140}{27/ \sqrt{157}} \\

= \frac{6}{2.1548} \\

= 2.78"

Decision Rule:

we reject H0 if Z>Z1-α or Z<-Z1-α

Z0.99= 2.33 (from table)

2.78>2.33

Reject H0

Hence, we can not conclude that the mean systolic blood pressure for a population of African-American is 140 mm Hg.